form an equivalent division problem for 5 divided by 1/3 by multiplying both the dividend and divisor by 3 then find the quotient

Answer by stanbon(75887)

form an equivalent division problem for 5 divided by 1/3 by multiplying both the dividend and divisor by 3 then find the quotient
-------------
x = 5/(1/3)
---
x = (5*3)/[(1/3)*3]
---
x = 15/1
---
x = 15
==========
Cheers,
Stan H.
==========